OpenStudy (anonymous):

What quandrant is 11pi/6 in? and how do you label the opposite, hypotenuse, and adjacent parts of the triangle? Why? please somebody help me

OpenStudy (anonymous):

\[\frac{11\pi}{6}\] is in quadrant 4

OpenStudy (anonymous):

look at the unit circle on the last page of the cheat sheet and you will see it

OpenStudy (anonymous):

 hypotenuse is 1, it is always 1. short side is\[ \frac{1}{2}\] long side is \[\frac{\sqrt{3}}{2}\]